OpenStudy (anonymous):

What is the sum of the rational expressions below? 3x-1/2x + 3x/x-1

OpenStudy (lgbasallote):

\[\large \frac{3x - 1}{2x} + \frac{3x}{x-1} = \frac{(3x-1)(x-1) + (2x)(3x)}{2x(x-1)}\] \[\large \frac{3x^2 - 4x + 1 + 6x^2}{2x(x-1)}\] \[\large \frac{9x^2 - 4x +1}{2x(x-1)}\]

OpenStudy (anonymous):

Take LCM first and then add both these The answer you may get is (9x^2-4x+1)/(2x^2-2x)

OpenStudy (anonymous):

@lgbasallote is right